OpenStudy (anonymous):

The mixed number -4/1/5 can be written as the improper fraction -19/5. true or false.

OpenStudy (misssunshinexxoxo):

What do you believe it is?

OpenStudy (anonymous):

false correct?

OpenStudy (texaschic101):

-4 1/5 disregard the negative sign for now....now take your whole number (4) and multiply it by the denominator (5), then add it back to the numerator (1)...and put that over the original denominator. Now add your negative sign. -4 1/5 = - (4 * 5 + 1) / 5 = - 21/5

OpenStudy (anonymous):

so is it false?

OpenStudy (texaschic101):

it is false